# Anomaly diagnosis via symmetry restriction in two-dimensional lattice systems

### v1

@@ 1 Introduction
Though they bear dynamical consequences, anomalies themselves are a purely kinematic property of the symmetry operators. For instance, in zero spatial dimensions (0D), [^1] Throughout, we denote $d$ spatial dimensions by $d$D. a unitary symmetry is anomalous when it acts on the Hilbert space as a nontrivial projective representation of the symmetry group $G$. This correspondence allows for a simple classification of 0D anomalies in terms of the cohomology group $H^{2}(G,U(1))$ [14, 15, 16]. [^2] We note that 0D anomalies are in one-to-one correspondence with 1D SPTs. In higher dimensions, we advocate that an anomaly should be viewed as a form of entanglement borne by the symmetry operators. For unitary, internal symmetries in 1D and 2D bosonic systems, this entanglement structure is characterized by indices valued in the cohomology groups $H^{3}(G,U(1))$ and $H^{4}(G,U(1))$, respectively [17, 18]. This scheme mirrors the classification of SPT phases in one higher dimension [11].
@@ 4 Obstruction to a trivially gapped Hamiltonian
In the previous section, we have defined a pair of indices, $[\nu]\in H^{2}(G,\mathbb{Q}_{+})$ and $[\omega]\in H^{4}(G,U(1))$, for any $G$-symmetry acting by FDQCs on a 2D lattice system. Here, we justify the interpretation of $[\omega]$ as a label for the anomaly of the symmetry. That is, we show—under certain physically reasonable assumptions—that a $G$-symmetry with nontrivial $[\omega]$ does not admit a $G$-symmetric invertible state. Thus, there cannot exist a symmetric gapped Hamiltonian with a unique invertible ground state.
@@
Our first assumption is that the defining procedure for the index $[\omega]$ can be appropriately modified to the setting of QCAs and FDQCs “with tails”. More precisely, a QCA with tails refers to a unitary that preserves locality of operators up to corrections that decay exponentially with distance. Similarly, a FDQC with tails refers to a circuit whose gates are supported on a finite number of sites, up to exponentially decaying corrections. [^16] We note that a FDQC with tails is equivalent to a locally generated unitary, i.e. finite time evolution by a time-dependent local Hamiltonian. In 1D, the GNVW index is well-defined in the setting of QCAs with tails, and the $\mathbb{Q}_{+}$ classification of 1D QCAs modulo FDQCs is unaffected by allowing for tails [37]. Thus, we anticipate that the procedure can be modified to allow for non-strict locality of all of the operators used to define $[\nu]$ and $[\omega]$. [^17] This would require a suitable notion of “restriction” of a QCA with tails. We emphasize that there are subtleties involving finite size effects arising from the lack of strict locality; we do not attempt a rigorous treatment here. Henceforth, the terms QCA and FDQC will refer to operators with tails.
@@
Bearing these subtleties in mind, we may now outline the argument. Recall that a state |$\ket{\psi}$ is invertible if there exists a state |$\ket{\phi}$ living in an ancillary Hilbert space, and a FDQC $V$ such that $V(\ket{\psi}\otimes\ket{\phi})$ is a product state [38, 39]. Further recall that a state |$\ket{\psi}$ is short-range entangled (SRE) if there exists a FDQC $W$ such that $W\ket{\psi}$ is a product state [40]. [^18] The standard definition of SRE states allows for tensoring with a product state in an ancillary Hilbert space. We will omit this possibility for simplicity; this does not affect the validity of Theorem 4.2. Note that in general, short-range entanglement is a strictly stronger condition than invertibility. [^19] We note that Kitaev uses an alternative definition of SRE that includes invertible states in the class of SRE states. Here, we use the stated definition. However, in 1D bosonic systems it is believed that all invertible states are also SRE [40]. [^20] In 1D fermionic systems there are invertible states that are not SRE, such as the Kitaev chain [41]. For our purpose we will assume this to be the case:
Assumption 4.1.
Every 1D invertible state is SRE.
Under this assumption, we demonstrate the following proposition:
@@
To begin, consider a restriction $U^{A}_{g}$ of $U_{g}$ to the disk $A$. By Lemma 4.3, there exists a FDQC $\Sigma_{g}$ supported near $\partial A$ such that $\Sigma_{g}U^{A}_{g}\ket{\psi}=\ket{\psi}$, for each $g\in G$. Let us redefine $U^{A}_{g}\to\Sigma_{g}U^{A}_{g}$, such that after the redefinition, $U^{A}_{g}\ket{\psi}=\ket{\psi}$. Since $\Omega_{g,h}=U^{A}_{g}U^{A}_{h}(U^{A}_{gh})^{-1}$, it follows that $\Omega_{g,h}\ket{\psi}=\ket{\psi}$.
@@
By Lemma 4.5, there exists a restriction $\Gamma^{a}_{g,h,k}$ of $\Gamma_{g,h,k}$ to the vicinity of $a$, such that $\Gamma^{a}_{g,h,k}\ket{\psi^{\prime}}=\ket{\psi^{\prime}}$. Altogether, by definition (29), it follows that $\omega(g,h,k,l)\ket{\psi^{\prime}}=\ket{\psi^{\prime}}$. Hence $\omega(g,h,k,l)=1$, and the anomaly index $[\omega]$ is trivial.
@@
Given a 2D SRE state |$\ket{\psi}$, a FDQC $U$ such that $U\ket{\psi}=\ket{\psi}$, and a restriction $U^{A}$ of $U$ to a disk $A$, there exists a FDQC $\Sigma$ supported near $\partial A$ such that $\Sigma U^{A}\ket{\psi}=\ket{\psi}$. [^21] We thank Michael Levin for insight on this point.
@@
The proof relies on Assumption 4.1. First, consider the case where |$\ket{\psi}$ is a product state |$\ket{\psi}=\bigotimes_{i}\ket{\psi}_{i}$.
@@
Therefore, |$\ket{\phi}_{\boldsymbol{\partial A}}$ is a 1D invertible state, hence it is SRE by Assumption 4.1. Consequently, there exists a FDQC unitary $\Sigma$, supported near $\boldsymbol{\partial A}$, such that $\Sigma\ket{\phi}_{\boldsymbol{\partial A}}=\ket{\psi}_{\boldsymbol{\partial A}}$. Since
@@ D.2 Ambiguity from $U_{g}^{A}$
To continue the procedure, we define the operator $\widetilde{T}_{g,h}=T_{\widetilde{\nu}(g,h)}^{-1}$ acting on $\mathcal{H}_{\partial}{A}$ [^28] In principle, the operator $\widetilde{T}_{g,h}$ may require ancilla qudits of dimension different from that of the original ancillary system. For simplicity we will assume that is not the case, and that $\widetilde{T}_{g,h}$ and $T_{g,h}$ act on the same Hilbert space $\mathcal{H}_{\partial A}$. Furthermore, let us define an operator $T_{g}=T_{\mu(g)}^{-1}$ for each $g\in G$, such that $\Sigma_{g}\otimes T_{g}$ is a 1D FDQC. Due to (15), we may decompose $\widetilde{T}_{g,h}$ as
@@
$=\bar{\Sigma}_{g}\cdot\hphantom{}^{g}\bar{\Sigma}_{h}\bar{\Omega}_{g,h}\bar{\Sigma}_{gh}^{-1}\cdot\hphantom{}^{\widetilde{gh}}\big{(}\Sigma^{L}_{k}\cdot\hphantom{}^{k}\Sigma^{L}_{l}\Omega^{L}_{k,l}(\Sigma^{L}_{kl})^{-1}\big{)}\bar{\Sigma}_{gh}\bar{\Omega}_{g,h}^{-1}\cdot\hphantom{}^{g}\bar{\Sigma}_{h}^{-1}\bar{\Sigma}_{g}^{-1}$ (122)

### v2

@@ 1 Introduction
Though they bear dynamical consequences, anomalies themselves are a purely kinematic property of the symmetry operators. For instance, in zero spatial dimensions (0D), [^1] Throughout, we denote $d$ spatial dimensions by $d$D. a unitary symmetry is anomalous when it acts on the Hilbert space as a nontrivial projective representation of the symmetry group $G$. This correspondence allows for a simple classification of 0D anomalies in terms of the cohomology group $H^{2}(G,U(1))$ [14, 15, 16]. [^2] We note that 0D anomalies are in one-to-one correspondence with 1D SPTs. In higher dimensions, we advocate that an anomaly should be viewed as a form of entanglement borne by the symmetry operators. For unitary, internal symmetries in 1D and 2D bosonic systems, this form of entanglement is characterized by indices valued in the cohomology groups $H^{3}(G,U(1))$ and $H^{4}(G,U(1))$, respectively [17, 18]. This scheme mirrors the classification of SPT phases in one higher dimension [11].
@@ 4 Obstruction to a trivially gapped Hamiltonian
In the previous section, we have defined a pair of indices, $[\nu]\in H^{2}(G,\mathbb{Q}_{+})$ and $[\omega]\in H^{4}(G,U(1))$, for any $G$-symmetry acting by FDQCs on a 2D lattice system. Here, we justify the interpretation of $[\omega]$ as a label for the anomaly of the symmetry. That is, we show—under certain assumptions—that a $G$-symmetry with nontrivial $[\omega]$ does not admit a $G$-symmetric invertible state. Thus, there cannot exist a symmetric, gapped, local Hamiltonian with a unique invertible ground state.
@@
We will work under the assumption that the defining procedure for the index $[\omega]$ can be appropriately modified to the setting of QCAs and FDQCs “with tails”. More precisely, a QCA with tails refers to a unitary that preserves locality of operators up to corrections that decay exponentially with distance. Similarly, a FDQC with tails refers to a circuit whose gates are supported on a finite number of sites, up to exponentially decaying corrections. [^16] We note that a FDQC with tails is equivalent to a locally generated unitary, i.e. finite time evolution by a time-dependent local Hamiltonian. In 1D, the GNVW index is well-defined in the setting of QCAs with tails, and the $\mathbb{Q}_{+}$ classification of 1D QCAs modulo FDQCs is unaffected by allowing for tails [37]. Thus, we anticipate that the procedure can be modified to allow for non-strict locality of all of the operators used to define $[\nu]$ and $[\omega]$. [^17] This would require a suitable notion of “restriction” of a QCA with tails. We emphasize that there are subtleties involving finite size effects arising from the lack of strict locality; we do not attempt a rigorous treatment here. Henceforth, the terms QCA and FDQC will refer to operators with tails.
@@
Bearing these subtleties in mind, we may now outline the argument. Recall that a state |$\ket{\psi}$ is invertible if there exists a state |$\ket{\phi}$ living in an ancillary Hilbert space, and a FDQC $V$ such that $V(\ket{\psi}\otimes\ket{\phi})$ is a product state [38, 39]. Further recall that a state |$\ket{\psi}$ is short-range entangled (SRE) if there exists a FDQC $W$ such that $W(\ket{\psi}\otimes\ket{\psi_{0}})$ is a product state, where |$\ket{\psi_{0}}$ is itself an ancillary product state [40]. Note that in general, short-range entanglement is a strictly stronger condition than invertibility. [^18] We note that Kitaev uses an alternative definition of SRE that includes invertible states in the class of SRE states. Here, we use the stated definition. However, in 1D bosonic systems it has been shown that all invertible states are also SRE [41]. [^19] In 1D fermionic systems there are invertible states that are not SRE, such as the Kitaev chain [42].
In the following proof, we will ignore the need to tensor with ancillas to trivialize SRE states. In other words, for any SRE state |$\ket{\psi}$, we assume there exists a FDQC $W$ such that $W\ket{\psi}$ is a product state. This assumption is not strictly necessary, and is merely invoked to streamline the proof.
We now demonstrate the following theorem:
@@
To begin, consider a restriction $U^{A}_{g}$ of $U_{g}$ to the disk $A$. By Lemma 4.2, for each $g\in G$ there exists a FDQC $\Sigma_{g}$ supported near $\partial A$ such that $\Sigma_{g}U^{A}_{g}\ket{\psi}=\ket{\psi}$. Let us redefine $U^{A}_{g}\to\Sigma_{g}U^{A}_{g}$, such that after the redefinition, $U^{A}_{g}\ket{\psi}=\ket{\psi}$. Since $\Omega_{g,h}=U^{A}_{g}U^{A}_{h}(U^{A}_{gh})^{-1}$, it follows that $\Omega_{g,h}\ket{\psi}=\ket{\psi}$.
@@
By Lemma 4.4, there exists a restriction $\Gamma^{a}_{g,h,k}$ of $\Gamma_{g,h,k}$ to the vicinity of $a$ such that $\Gamma^{a}_{g,h,k}\ket{\psi^{\prime}}=\ket{\psi^{\prime}}$. Altogether, by definition (29), it follows that $\omega(g,h,k,l)\ket{\psi^{\prime}}=\ket{\psi^{\prime}}$. Hence $\omega(g,h,k,l)=1$, and the anomaly index $[\omega]$ is trivial.
@@
Given a 2D SRE state |$\ket{\psi}$, a FDQC $U$ such that $U\ket{\psi}=\ket{\psi}$, and a restriction $U^{A}$ of $U$ to a disk $A$, there exists a FDQC $\Sigma$ supported near $\partial A$ such that $\Sigma U^{A}\ket{\psi}=\ket{\psi}$. [^20] Again, we will ignore the need to tensor with ancillas to trivialize SRE states., [^21] We thank Michael Levin for insight on the proof of this lemma.
@@
This proof uses the fact that every 1D invertible state is SRE, as shown in [41]. First, consider the case where |$\ket{\psi}$ is a product state |$\ket{\psi}=\bigotimes_{i}\ket{\psi}_{i}$.
@@
Therefore, |$\ket{\phi}_{\boldsymbol{\partial A}}$ is a 1D invertible state, hence it is SRE. Consequently, there exists a FDQC unitary $\Sigma$, supported near $\boldsymbol{\partial A}$, such that $\Sigma\ket{\phi}_{\boldsymbol{\partial A}}=\ket{\psi}_{\boldsymbol{\partial A}}$. Since
@@ D.2 Ambiguity from $U_{g}^{A}$
To continue the procedure, we define the operator $\widetilde{T}_{g,h}=T_{\widetilde{\nu}(g,h)}^{-1}$ acting on $\mathcal{H}_{\partial{A}}$. [^28] In principle, the operator $\widetilde{T}_{g,h}$ may require ancilla qudits of dimension different from that of the original ancillary system. For simplicity we will assume that is not the case, and that $\widetilde{T}_{g,h}$ and $T_{g,h}$ act on the same Hilbert space $\mathcal{H}_{\partial A}$. Furthermore, let us define an operator $T_{g}=T_{\mu(g)}^{-1}$ for each $g\in G$, such that $\Sigma_{g}\otimes T_{g}$ is a 1D FDQC. Due to (15), we may decompose $\widetilde{T}_{g,h}$ as
@@
$=\bar{\Sigma}_{g}\cdot\hphantom{}^{g}\bar{\Sigma}_{h}\bar{\Omega}_{g,h}\bar{\Sigma}_{gh}^{-1}\cdot\hphantom{}^{\widetilde{gh}}\big{(}\Sigma^{L}_{k}\cdot\hphantom{}^{k}\Sigma^{L}_{l}\Omega^{L}_{k,l}(\Sigma^{L}_{kl})^{-1}\big{)}\bar{\Sigma}_{gh}\bar{\Omega}_{g,h}^{-1}\cdot\hphantom{}^{g}\bar{\Sigma}_{h}^{-1}\bar{\Sigma}_{g}^{-1}\left(\hphantom{}{}^{\widetilde{g}\cdot\widetilde{h}}\widetilde{\Omega}^{L}_{k,l}\right)^{-1}.$